OpenStudy (anonymous):

Factor completely: 2x^3 – 18x

OpenStudy (jdoe0001):

take common factors firstly then recall that \(\bf a^2-b^2 = (a-b)(a+b)\)

OpenStudy (anonymous):

these are the answers so im confuzed 2x(x – 9) x(x – 3)(x – 3) 2x(x + 3)(x – 3) 2(x + 3)(x – 3)

OpenStudy (jdoe0001):

so if you take common factor to \(\bf 2x^3 – 18x\) what would you get?

OpenStudy (anonymous):

umm...

OpenStudy (jdoe0001):

hmmm *scratch chin * hmm what would be the common factor....

OpenStudy (anonymous):

2

OpenStudy (jdoe0001):

just 2? anything else common to both terms?

OpenStudy (anonymous):

i dont think so

OpenStudy (jdoe0001):

well, let's take a peek -> \(\bf 2x^3 - 18x\implies {\color{blue}{ 2}}x^3 - ({\color{blue}{ 2}}\cdot 9)x\implies {\color{blue}{ 2}}\cdot {\color{blue}{ x}}\cdot x\cdot x - ({\color{blue}{ 2}}\cdot 9){\color{blue}{ x}}\) what do you think?

OpenStudy (anonymous):

uh...